question a i ii iii16. ChargeBank Plc is a battery manufacturer with securities listed on a large exchange. The company has the following capita
Answer #1

i) Caculation of Current Share price Using Dividend Growth model Current Share Price = D1/ K-g Caculation of k using CAPM mod

i) Caculation of Current Share price Using Dividend Growth model Current Share Price = D1/ K-g Caculation of k using CAPM model k = Risk free rate+ beta ( Market return - Risk Free rate) 0.03+1.2*(.08-.03) 9% Given Divivdend = .65 Growth rate = 4% Current Share Price = D1/ K-g ='0.65/0.09-0.04 £ 13.00 Therefore current share Price = £ 13.00 ii) Caculation of after tax cost of Debt Unsecured Bond Cost of Debt unsecured before tax = 5.75% Tax rate = 20% Cost of Debt unsecured after tax = 5.75%-(5.75%*20%) 4.60% iii) Caculation of Current Weighted Average Cost of Capital WACC = ke* Equity/(Equity+Debt) +kd(1-t) *Debt/Equity+Debt Market value of common stock = Outstanding shares * Market share = 12000000*£ 13 £ 15,60,00,000.00 Market value of Debt = Outstanding Bond * market value (150000000/100)*90.12 £ 13,51,80,000.00 Bank Loan = £ 55000000 Total Funds = 156000000+135180000+55000000 £34,61,80,000.00 WACC = 0.09*(156000000/346180000) +0.05*(135180000/346180000) +0.06(1-0.2)*(55000000/346180000) 0.07 6.77%